OpenStudy (anonymous):

If f(x) = x − 4 and g(x) = 1 over 2x − 2, find f[g(−6)].

OpenStudy (anonymous):

9 3 −9 −3

OpenStudy (anonymous):

What is \(g(-6)\)? Can you at least find that first?

OpenStudy (anonymous):

\[f(x)=x-4\]\[g(x)=\frac{ 1 }{ 2x-2 }\]Then,\[f(g(x))=\frac{ 1 }{2x-2 }-4\]Then,\[f(g(-6))=\frac{ 1 }{ 2(-6)-2 }-4\]\[f(g(-6))=\frac{ 1 }{ -14 }-\frac{ 4 }{ 1 }\]Subtracting both fractions,\[f(g(-6))=\frac{ 1-(4)(-14) }{ -14 }=-\frac{ 57 }{ 14 }\]

OpenStudy (anonymous):

ok I see

OpenStudy (anonymous):

wait is it c

OpenStudy (anonymous):

\[-\frac{ 57 }{ 14 } \approx -4.07\]Either I did something wrong, the functions you gave were wrong, or the answers they gave you are wrong.

OpenStudy (anonymous):

this is right its in a book I have to do